A golf course sprinkler system discharges water from a horizontal pipe at the rate of 7500 cm3/s . At one point in the pipe, where the radius is 4.05 cm , the water's absolute pressure is 2.25×105 Pa . At a second point in the pipe, the water passes through a constriction where the radius is 2.20 cm .

What is the water's absolute pressure as it flows through this constriction?​

P= ___________Pa
Answer #1

Solution,

Using Bernoulli's equation,

P1+ 0.5 pv1^2 = P2 +0.5 pv2^2

Discharge, Q = 7500 cm^3/sec

(pi r1^2) v1 = 7500

Velocity, v1 = 145.55 cm/s = 1.45 m/s

At other end, Q = 7500

(pi x r2^2) v2 = 7500

Velocity, v2 = 493.25 cm/s = 4.93 m/s

Substituting in equation,

(2.25 x 10^5) + 0.5(1000 x 1.45^2) = P2 + 0.5(1000 x 4.93^2)

Pressure, P2 = 2.14 x 10^5 Pa

Comment in case any doubt please rate my answer . ...